OpenStudy (anonymous):

The measure of an angle is 78° less than the measure of its complement. What is the measure of the angle?

OpenStudy (anonymous):

The measure of an angle is 78° less than the measure of its complement. What is the measure of the angle?

OpenStudy (anonymous):

The angle and its complement add up to 90 degrees, right? That's what it means to be a complementary angle. So you need two angle measures that add to 90, and one number has to be 78 larger than the other.

OpenStudy (anonymous):

Besides knowing what complementary means, this is mostly an algebra word problem.

OpenStudy (anonymous):

If 12 is one angle, then the other angle would be 78, since they add to 90. But 12 is not 78 less than 78, so "12" doesn't satisfy both requirements.

OpenStudy (anonymous):

is 12 or 51 or 84 or 6

OpenStudy (anonymous):

Try 6: If it's 6, then the complement is what you have to add to get 90... so the complement is 84. Then check: if it's 6, and the complement is 84, then is 6 equal to 78 less than 84? Yes. So it's 6...

OpenStudy (anonymous):

Read the following two statements. Then, if possible, use the Law of Detachment to draw a conclusion. If x = 3, then 3x – 4 = 5. x = 3 .If 3x – 4 = 5 then x = 3 3x – 4 = 5 x = 3 not possible

OpenStudy (anonymous):

I don't remember what law of detachment means. It looks like they are saying "If x = 3 then you also know that "3x - 4 = 5" is true" So then the 2nd statement says the x = 3 (like it gives you that as "FACT") So it would seem to me that you would conclude that the 2nd part of the first statement is true... Not 100% sure though...

OpenStudy (anonymous):

Name the property of equality that justifies the following statement If p = q, then p-r=q-r Multiplication Property Subtraction Property Symmetric Property Reflexive Property